OpenStudy (is3535):

A parallelogram has an area of 224 square centimeters and a height of 14 centimeters. What is the length of the base of the parallelogram?

OpenStudy (is3535):

@butterflydreamer

OpenStudy (butterflydreamer):

area of parallelogram = base * height So using the info we're given. 224 = base * 14 To find base, divide both sides by 14 :) base = 224/14 = ?

OpenStudy (is3535):

14

OpenStudy (is3535):

sryy 16

OpenStudy (butterflydreamer):

yep ^_^ 16 is correct